### v1

# WaterVG: Waterway Visual Grounding based on Text-Guided Vision and mmWave Radar

Abstract.
The perception of waterways based on human intent holds significant importance for autonomous navigation and operations of Unmanned Surface Vehicles (USVs) in water environments. Inspired by visual grounding, in this paper, we introduce WaterVG, the first visual grounding dataset designed for USV-based waterway perception based on human intention prompts. WaterVG encompasses prompts describing multiple targets, with annotations at the instance level including bounding boxes and masks. Notably, WaterVG includes 11,568 samples with 34,950 referred targets, which integrates both visual and radar characteristics captured by monocular camera and millimeter-wave (mmWave) radar, enabling a finer granularity of text prompts. Furthermore, we propose a novel multi-modal visual grounding model, Potamoi, which is a multi-modal and multi-task model based on the one-stage paradigm with a designed Phased Heterogeneous Modality Fusion (PHMF) structure, including Adaptive Radar Weighting (ARW) and Multi-Head Slim Cross Attention (MHSCA). In specific, MHSCA is a low-cost and efficient fusion module with a remarkably small parameter count and FLOPs, elegantly aligning and fusing scenario context information captured by two sensors with linguistic features, which can effectively address tasks of referring expression comprehension and segmentation based on fine-grained prompts. Comprehensive experiments and evaluations have been conducted on WaterVG, where our Potamoi archives state-of-the-art performances compared with counterparts.
Visual grounding, Multi-modal fusion, USV-based water-surface perception, mmWave radar
This work is partially supported by the XJTLU AI University Research Centre and Jiangsu Province Engineering Research Centre of Data Science and Cognitive Computation at XJTLU. Also, it is partially funded by the Suzhou Municipal Key Laboratory for Intelligent Virtual Engineering (SZS2022004) as well as funding: XJTLU-REF-21-01-002, XJTLU- RDF-22-01-062, and XJTLU Key Program Special Fund (KSF-A-17). This work received financial support from Jiangsu Industrial Technology Research Institute (JITRI) and Wuxi National Hi-Tech District (WND).Runwei Guan, Liye Jia and Fengyufan Yang contribute equally to this work.*Corresponding email: yutaoyue@hkust-gz.edu.cn.
Figure 1. Overview of our pipeline for WaterVG. The camera and radar align prompt on appearance, motion and distance features.
1. Introduction
With the rapid advancement of deep learning technologies, intelligent perception in water-surface environments has witnessed substantial progress in various domains (Yao et al., 2023a, b), including autonomous driving of USVs (Yao et al., 2023c), maritime intelligent transportation (Guo et al., 2023), etc. Specifically, current advancements in water-surface intelligent perception primarily encompass instance-level detection (Cheng et al., 2021c), segmentation(Guan et al., 2023c, b), and tracking of waterborne targets (Zhang et al., 2021), detection of drivable areas (Guan et al., 2024), waterline detection (Guan et al., 2023d), Simultaneous Localization And Mapping (SLAM) of waterways (Cheng et al., 2021a), and related facets. Based on the aforementioned investigation, our findings reveal that current research predominantly focuses on sensor-based automatic perception of water surface environments.
However, there is a paucity of research on human-guided specific target perception through natural language, which belongs to the domain of visual grounding. In the context of visual grounding in waterways, humans can convey target appearance, position, distance and related features through descriptions, facilitating more intuitive and efficient guidance for USVs. This leads to enhanced precision in navigation and localization, thereby holding significant implications for waterway intelligent traffic systems, search and rescue operations, and aquatic environment conservation. Consequently, our focus is centered on the domain of visual grounding, encompassing two primary tasks: Referring Expression Comprehension (REC) (Qiao et al., 2020) and Referring Expression Segmentation (RES) (Chrupała, 2022). These tasks involve presenting a prompt that describes one or more targets, and the visual grounding model is tasked with predicting bounding boxes and masks of the objects in the image that most closely match the description provided in the prompt.
Building upon the aforementioned considerations, we contemplate the establishment of an optimal visual grounding dataset tailored for robust perception and monitoring in waterway environments. Firstly, pure visual sensors are susceptible to interference in adverse weather and lighting (Cheng et al., 2021c; Yao et al., 2023a, c). Secondly, visual sensors can only capture qualitative features of targets in terms of appearance or position. but traffic supervisors often need quantitative distance and motion information about waterway targets. Therefore, when one desires to ascertain whether there is a vessel approaching behind the fog or whether obstacles exist within a specific distance, visual sensors are unable to precisely locate targets based on human intent. To enable target referring flexibly, we introduce 4D mmWave radar to complement the visual sensor, which as an all-weather perception sensor, can detect targets in the form of point clouds, offering higher resolution compared to conventional mmWave radars (Guan et al., 2023b, d; Cheng et al., 2021b). Concurrently, while visual inputs provide appearance features, including color, size, texture and shape, radar can furnish target reflection characteristics (Yao et al., 2023b), including distance, velocity, reflected power, etc.
Figure 2. Samples in WaterVG, containing referred targets with prompts under various situations. The annotations include bounding boxes and masks, supplemented by radar point clouds (red).
Upon these insights, we develop the first visual grounding dataset for water-surface perception, WaterVG, based on the WaterScenes dataset (Yao et al., 2023c), which is a large-scale USV-based waterway perception dataset integrating the monocular camera and 4D mmWave radar. Detailedly, diverging from other vision-based visual grounding datasets, our WaterVG dataset has the following distinguishing features: i) Manual fine-grained prompts with a maximum of 40 tokens to describe targets not only with visual attributes like color, size, shape and texture but also incorporate features imperceptible to vision yet detectable by radar, such as numeric distance and motion characteristics. ii) Flexible referred targets with 1 to 15 per prompt. iii) Diverse prompts, containing specific descriptions, coarse and fine queries about one or more of the target’s characteristics.
Correspondingly, we propose a novel baseline model, Potamoi, adept at seamlessly integrating these three distinctly heterogeneous modalities. The crux of fusing heterogeneous modalities lies in how to align their diverse features within a latent space. Building on this, we propose an effective fusion structure called Phased Heterogeneous Modality Fusion (PHMF). PHMF performs a phased fusion of visual, radar, and linguistic features. This approach not only effectively alleviates the complexity increase and speed reduction caused by the superposition of multiple modalities with multiple branching channels but also deepens the neural network, enhancing the model’s expressive capacity. PHMF’s core lies in our proposed multi-modal fusion method called Multi-Head Slim Cross Attention (MHSCA). This module within Potamoi efficiently fuses abundant contextual information captured by two sensors with fine-grained linguistic features, achieving exemplary performance on both REC and RES tasks with lower parameter numbers and FLOPs than vanilla Multi-Head Cross Attention (MHCA) (Wu et al., 2023) and Multi-Head Linear Cross Attention (MHLCA) (Choromanski et al., 2020).
In summary, the contributions of this paper are as follows:
  1. (1)
    We establish the first visual grounding dataset for USV-based waterway perception, WaterVG, which comprises 11,568 triplet input data, including images, radar point clouds, and prompts. Moreover, the ground truth encompasses two types of instance-level annotations: bounding boxes and masks. In particular, our textual prompts contain extra corresponding descriptions of target distance and motion information.
  2. (2)
    We propose a novel one-stage multi-task baseline model, Potamoi, where an effective fusion network structure called Phased Heterogeneous Modality Fusion (PHMF) aligns and fuses three modalities. PHMF includes Adaptive Radar Weighting (ARW) modules and a low-cost fusion method called Multi-Head Slim Cross Attention (MHSCA). MHSCA can fuse heterogeneous modalities efficiently for visual grounding with few parameters, low FLOPs and power than vanilla Multi-Head Cross Attention (MHCA) (Dao et al., 2022) and Multi-Head Linear Cross Attention (MHLCA) (Choromanski et al., 2020).
  3. (3)
    Comprehensive experiments based on the WaterVG dataset encompasses various aspects such as the performance of visual grounding models under fusion and vision patterns, and multi-modal fusion methods. These experiments provide further insights and deepen our understanding of the field.
2. Related Works
2.1. Datasets of Visual Grounding
As a challenging branch within multi-modal learning, visual grounding aims to locate, based on natural language query, the targets in images that best match the descriptions provided by the natural language (Zhou et al., 2023). This is further divided into REC and RES tasks. The REC task focuses on localizing targets in the form of bounding boxes, whereas RES aims to delineate target masks. The majority of current 2D visual grounding datasets (Kazemzadeh et al., 2014; Plummer et al., 2015; Yu et al., 2016; Krishna et al., 2017; Deruyttere et al., 2019; Wang et al., 2020; Chen et al., 2020; Zhan et al., 2023; Liu et al., 2023; Wu et al., 2023; He et al., 2023) primarily cater to REC tasks, relying solely on pure visual data, lacking other modalities queried by prompts. Moreover, there is a current trend towards a prompt being capable of querying multiple targets, deviating from the single-target prompt approach as seen in datasets like RefCOCOs (Yu et al., 2016), thereby aligning more flexibly with real-world scenarios. Additionally, multi-task learning stands as a significant research focus within visual grounding. In contrast to other datasets, our WaterVG supports both REC and RES tasks, accommodating natural language reference to multiple targets and facilitating multi-task learning. Notably, we provide image and radar point cloud data concurrently, enabling more precise queries based on natural language prompts.
2.2. Visual Grounding Models
As visual grounding advances, its models are primarily categorized into three major classes: two-stage (Liu et al., 2019; Yang et al., 2019b; Yu et al., 2018; Wang et al., 2019; Hong et al., 2019), one-stage (Luo et al., 2020; Chen et al., 2018; Yang et al., 2019a), and transformer-based models (Zhu et al., 2022; Du et al., 2022; Deng et al., 2021). Among them, firstly, two-stage models split the generation of Regions Of Interest (ROI) and target classification into two parts, where the first stage is responsible for matching the regions of interest with text features while the second stage is to classify the foreground and background categories of ROI based on the fused features of text and image. Therefore, two-stage models usually adopt Faster R-CNN (Girshick, 2015) as the skeleton. Secondly, one-stage models usually combine bounding box regression and classification into one step. Besides, linguistic features usually fuse with image features at the end of the image backbone or within Feature Pyramid Networks (FPN). Thirdly, transformer-based models are usually in the format of an image backbone and a text encoder followed by a transformer-based encoder and decoder, where the fusion operation happens before the transformer-based encoder. There are many different text-visual fusion methods among the aforementioned models, including cross-modal spatial attention (Zhu et al., 2022), multi-head cross attention (Wu et al., 2023), channel concatenation (Deng et al., 2021), and so on.
3. The WaterVG Dataset
Figure 3. Five-step prompt annotations based on the samples of image and projected radar point clouds.
Overall, the WaterVG dataset comprises 11,568 samples, including 9,254 in the training set, 578 in the validation set, and 1,736 in the test set. Each sample consists of an image, a frame of 4D radar point cloud, a prompt description, and ground truth presented in the forms of bounding boxes and masks. This section will sequentially elaborate on the construction process (Section 3.1) and data statistics (Section 3.2) of the WaterVG dataset.
3.1. Dataset Construction
Our WaterVG dataset is built upon the WaterScenes dataset (Yao et al., 2023c), a high-quality, multi-sensor dataset with vision and radar, specifically designed for various water-surface environments. It features high-resolution images from a monocular camera with 1920 × 1080 pixels, complemented by 4D radar point clouds, which are calibrated and time-synchronized with the monocular camera. Additionally, WaterScenes (Yao et al., 2023c) provides high-quality annotations for object detection in the form of bounding boxes, as well as instance segmentation masks. Building upon these resources, we engage two experts with extensive experience in intelligent aquatic transportation to conduct prompt annotation of our WaterVG dataset, which selects 11,568 representative samples from WaterScenes under different scenarios and lighting conditions. As illustrated in Fig. 3, the annotation process is divided into five distinct stages.
(1) Select candidate targets. To facilitate efficient annotation by experts, we develop an annotation GUI client specifically for the WaterVG dataset. In this system, the radar point clouds are projected onto the camera plane using the projection matrix based on the exterior parameters of the radar and camera, as well as the intrinsic parameters of the camera, with the point clouds displayed in red. Each point cloud is accompanied by information displaying the distance of the target from the USV and its radial velocity. Correspondingly, both bounding box and mask annotations are mapped onto the image. Annotators simply need to click on the location of the target to display these two types of annotations.
(2) Check the annotations. Once the bounding box and mask of a clicked object are displayed on the image, two annotators proceed to check the corresponding ground truth for accuracy. If the ground truth is verified as correct, they continue the next step. If not, the target is re-annotated to ensure the precision of the dataset.
(3) Write referring prompts. Based on the images and radar point clouds, the two annotators include as many physical characteristics of the targets as possible in the annotation prompt. These encompass the target’s category, size, color, and shape features perceptible by visual means. Furthermore, if the target is detected by the radar precisely accompanied by dense projected point clouds, we use characteristics of point clouds to further describe the target’s distance and motion trends from our USV. Additionally, if multiple targets are involved, annotators will also work on expressing their spatial location relationships. Based on these, the annotators aim to diversify the prompts by incorporating various sentence structures and methods of referencing the target.
(4) Enrich prompts. When facing identical scenes or targets in different samples, the limited vocabulary and subjective tone of the annotators are unavoidable. To enhance the diversity of prompt content, we utilized the existing annotated prompts and fed them into the ChatGPT-3.5-Turbo API, aiming to enrich and diversify the prompts by generating semantically equivalent yet more elaborate and varied expressions.
(5) Check each other’s prompt annotations. To avoid ambiguity in text prompts, as a final step, we exchange prompts written by two experts. Each expert reviews and examines the text prompt written by the other, respecting their respective writing styles. In cases where there is ambiguity or misidentification in target references, the two experts engage in discussions and reach a consensus for necessary modifications.
Figure 4. The statistics of WaterVG. (a) waterway proportion, (b) sentence pattern proportion, (c) query type proportion, (d) prompt length per sample, (e) word cloud, (f) co-occurrence of categories, (g) referred target number per sample, (h) category distribution of referred targets.
3.2. Dataset Statistics
Figure 4 presents the statistics of WaterVG from various perspectives. Figure 4(a) shows the waterway scenario proportion in WaterVG, including canal, river, lake and moat. Figure 4(b) presents the proportion of sentence patterns for prompts, including statement (egocentric), question, statement (imperative) and exclamation. Figure 4(c) is the prompt type proportion, containing 4 main types of query by exact features, query by partial features, query by number and query by category. Figure 4(d) illustrates the distribution of text prompt lengths, ranging from 2 to 40, encompassing both simple phrases and detailed long sentences. The word cloud in Figure 4(e) of WaterVG displays high-frequency words, including categories, appearance, location, distance, motion, spatial relationships, etc., reflecting the richness and comprehensiveness of the dataset text prompt. Figure 4(f) shows the co-occurrence matrix of referred targets for various categories, which partly reflects the category diversity of the targets described in each of our samples. Figure 4(g) illustrates the distribution statistics of referred target number per sample. In WaterVG, each text prompt refers to 1-15 targets, with the majority consisting of 1-2 targets. Figure 4(h) presents the distribution of referred target categories, including piers, buoys, sailors, ships, boats, pleasure boats and kayaks, a total of 34,950 referred targets. The above demonstrates the diversity and comprehensiveness of our WaterVG dataset in terms of visual grounding in waterway navigation. This is primarily reflected in the diversity of waterway scenes, richness of textual prompts in terms of content, patterns, and types, diversity of mentioned target categories, and flexibility in the number of mentioned targets.
Figure 5. The architecture of Potamoi. The text encoder (ALBERT) (Lan et al., 2020) for text prompts is frozen during training.
4. Methods
We propose a one-stage multi-task visual grounding network called Potamoi to simultaneously conduct two grounding tasks, REC and RES. We first introduce the overall pipeline of Potamoi as Figure 5 shows, and then illustrate our proposed Phased Heterogeneous Modality Fusion (PHMF), including (a) Adaptive Radar Weighting (ARW) and (b) Multi-Head Slim Cross Attention (MHSCA) in detail.
4.1. Overall Pipeline
Given an input triplet of an image ($I\in R^{3\times I_{H}\times I_{W}}$), a frame of radar point clouds ($R$) and a text prompt $T$, feature encoders for three modalities are for mapping them to high-dimension latent space. Firstly, the image encoder adopts FastViT-SA12 (Vasu et al., 2023) with the cascade of reparametrized RepMixer and Multi-Head Self-Attention (MHSA). Secondly, for 3D radar point clouds, we primarily project point clouds from radar coordinate to image coordinate, and obtain a three-channel radar point cloud feature map called RVP maps. Each point cloud contains the target Range (R), compensated Velocity (V) and reflected Power (P). For the RVP map with sparse point cloud features, we adopt the combination of convolution, batch normalization and ReLU activation as the basic block. Each stage of radar backbone contains three basic blocks, where the convolution kernels are $3\times 3$, $1\times 1$ and $3\times 3$ with the stride of $1$, $1$ and $2$ for local feature extraction, feedforward and downsampling. Whatever image or radar encoder, there are four stages and one stem layer, and the stem layer downsamples the spatial size four times. The feature dimension of stage $i\in\{1,2,3,4\}$ is $(64\times 2^{i-1})\times\frac{I_{H}}{2^{i+1}}\times\frac{I_{W}}{2^{i+1}}$.
Importantly, the last three stage features from the radar encoder go through three ARW modules and fuse with the same-dimensional feature maps from the image encoder by element-wise addition in the vision-radar fusion network, whose outputs are three-stage sensor features. These sensor features then act as queries to fuse with the keys and values of textual features from the pretrained ALBERT (Lan et al., 2020) by our proposed MHSCA and output three-stage fusion features. Then, a Path Aggregation Network (PANet) (Li et al., 2022) is to fuse features with various receptive fields, and outputs four-stage pyramid features from $P2$ to $P5$.
We feed all four pyramid features to REC heads, which are transplanted and adapted from YOLOv8 decouple head (Jocher et al., 2023). Each pyramid feature corresponds to one REC head, containing a confidence branch and a bounding box regression branch. The shape of REC output for the stage $i\in\{2,3,4,5\}$ is $(1+4\times RegMax)\times\frac{I_{H}}{2^{i+1}}\times\frac{I_{W}}{2^{i+1}}$, where $1$ is the confidence score and $4$ is four coordinates of bounding box. $RegMax$ is the hyperparameter of Distributed Focal Loss (DFL) (Li et al., 2020). Furthermore, following the paradigm of PP-LiteSeg (Peng et al., 2022), we feed $P3$, $P4$ and $P5$ upon SPPM to RES head, which is based on Flexible Lightweight Decoder (FLD) (Peng et al., 2022). The output shape of RES head is $2\times I_{H}\times I_{W}$, including referred foreground and unreferred background categories.
4.2. Phased Heterogeneous Modality Fusion
We adopt a divide-and-conquer strategy for three modalities, including Adaptive Radar Weighting and Multi-Head Slim Cross Attention, which are vision-radar and text-sensor fusion in essence.
Figure 6. The structure of Adaptive Radar Weighting.
(a) Adaptive Radar Weighting. Firstly, radar, as an auxiliary modality, can provide target prior information on distance, speed (motion), and target reflectivity to assist in distinguishing target categories. When combined with images for text queries, two considerations arise regarding the RVP map. Firstly, not all text prompts contain distance or motion information, corresponding to the channel features of the RVP map. Secondly, radar can be affected by clutter and multi-path effects, leading to generation of useless clutter point clouds in environments, corresponding to the spatial features of the RVP map. To adaptively balance the role of RVP map in visual grounding, we propose a module called Adaptive Radar Weighting (ARW) to adaptively weigh the RVP map. As Figure 6 presents, assuming the radar feature is $F_{R}\in R^{C\times H\times W}$, we exert a linear projection with GeLU activation on it to obtain $\hat{F_{R}}\in R^{C\times H\times W}$. The resulting feature map is then prepared for adaptive channel attention matrix $W_{CA}\in R^{C}$ by applying adaptive average pooling to collapse spatial dimensions and computing attention weights by softmax. $\hat{F_{R}}$ is position-wise multiplication with $W_{CA}$ to obtain the channel-weighting feature $F_{R}^{CW}\in R^{C\times H\times W}$. Then spatial weighting matrix $W_{SA}\in R^{1\times H\times W}$ is achieved by convolving $F_{R}^{CW}$ with a GELU activation and a 1 $\times$ 1 convolution, followed by element-wise multiplication with the original features and obtain the final radar feature $F_{R}^{ARW}\in R^{C\times H\times W}$ with both channel and spatial adaptive weighting. Equation 1 and 2 present the whole process above.
[…]
Figure 7. The structure of Slim Cross Attention.
(b) Multi-Head Slim Cross Attention. After obtaining three-stage sensor features ($F^{2}_{S}$, $F^{3}_{S}$ and $F^{4}_{S}$) containing the contextual features of vision and prior target features captured by radar, there are three linear projection branches from ALBERT to align the channel dimensions of three-stage sensor features. To effectively fuse two heterogeneous modalities with a relatively low cost, we propose Multi-Head Slim Cross Attention (MHSCA) with linear complexity. Given a sensor feature $F_{S}\in R^{C\times H\times W}$ and a text feature $F_{T}\in R^{L\times C}$, we first exert learnable encodings (Dosovitskiy et al., 2020)(Vaswani et al., 2017), $P_{I}$ and $P_{T}$ on them, respectively. Then we adopt 2D depth-wise convolution with $1\times 1$ kernel ($W_{Q}$) to obtain the sensor query ($Q$) while two 1D depth-wise convolutions ($W_{K}$ and $W_{V}$) are for generate key ($K$) and value ($V$). Equation 3 presents the aforementioned process.
$Q=F_{S}W_{Q}+P_{S},K=F_{T}W_{K}+P_{T},V=F_{V}W_{V}+P_{T},$
(3) $Q\in R^{C\times H\times W},K\in R^{L\times C},V\in R^{L\times C},$
Further, we employ adaptive pooling to adaptively downsample the sensor query $Q$ (Han et al., 2023), preserving its essential features while reducing its spatial size to obtain the sensor agent $A\in R^{C\times h\times w}$. $A$, as the agent of $Q$, is firstly flatten along spatial dimension as $A\in R^{n\times C}$ and conducts softmax attention ($\sigma(\cdot)$) with $K$ and $V$ to calculate the agent-conditional Value matrix $V_{A}$. (Equation 4)
$A=Flat(A)^{T},A\in R^{n(h\times w)\times C},$
(4) $V_{A}=\sigma(A,K,V)=\frac{AK^{T}}{\sqrt{d}}\cdot V,$
where $d$ is the channel dimension of $F_{S}$ and $F_{T}$.
Then, $V_{A}$ is forwarded to the second softmax attention, broadcasting the inter-modal similarity embedded in text feature to the sensor query $Q$, where $A$ acts as the key and $V_{A}$ is the value (Equation 5). Here $Q$ is also firstly flattened as $Q\in R^{N\times C}$, where $N>n$.
$Q=Flat(Q)^{T},Q\in R^{N(H\times W)\times C},$
(5) $\sigma(Q,A,V_{A})=\frac{QA^{T}}{\sqrt{d}}\cdot V_{A},$
After that, a residual path from the sensor feature $F_{S}$ is added to the output of the second softmax attention and obtains the image-like sensor-conditional text features $F_{TS}$ (Equation 6). Finally, a 2D depth-wise convolution with $1\times 1$ kernel size ($W$) is exerted on $F_{TS}$ to further feedforward.
(6) $F_{TS}=(Reshape(\sigma(Q,A,V_{A}))+F_{S})W,F_{TS}\in R^{C\times H\times W},$
Here, the computation complexity of our proposed slim cross attention is $O(Nnd)+O(NLd)$, which is lower than vanilla cross attention’s (Wu et al., 2023) $O(NLd)+O(NLd)$, as $n\ll L$ in practice.
4.3. Training Objectives
We reformulate the REC and RES tasks as the bounding box regression and semantic segmentation with one category. For both two tasks, the area occupied by the referred objects in the image is much smaller than the background area. Equation 7 presents the REC loss. We introduce Distributed Focal Loss (DFL) (Li et al., 2020) to alleviate the unbalanced distribution between two areas. Meanwhile, CIoU (Zheng et al., 2020) is added to refinedly regress the bounding box size by taking its aspect ratio into account.
(7) $L_{REC}=L_{DFL}+L_{CIoU},$
For RES task, we combine the dice loss with focal loss to jointly subtly optimize the mask prediction.
(8) $L_{RES}=L_{Dice}+L_{FL},$
We empirically argue that in the joint optimization process of REC and RES, the loss functions are not in the same order of magnitude. To avoid the domination of optimization by the task with a large loss, we introduce uncertainty weighting (Kendall et al., 2018) to balance multi-task optimization. Equation 9 demonstrates the loss function during the joint optimization of REC and RES.
(9) $L_{Joint}=\frac{1}{2\sigma_{1}^{2}}L_{REC}+\frac{1}{2\sigma_{2}^{2}}L_{RES}+log\sigma_{1}+log\sigma_{2},$
where $\sigma_{1}$ and $\sigma_{2}$ are learnable parameters to weigh two loss items.
Figure 8. Images with projected radar point clouds (PCs) (first row) and prediction results by Potamoi (second row) and MCN (Luo et al., 2020) (third row) under different situations. Red bounding boxes are REC predictions by text prompt while translucent red masks are RES predictions. Blue PCs denote clutter while other represent targets. The small gray letter on the PC refers to the distance (m) from targets to the ego USV.
5. Experiments
5.1. Settings of Dataset and Implementation
Datasets. Our proposed WaterVG is for training (9,254), validation (578) and testing (1,736) on all models. Besides, our WaterVG includes two benchmarks, Fusion-based Visual Grounding (FVG) and Vision-based Visual Grounding (VVG).
Models. For the REC, MattNet (Yu et al., 2018), MCN (Luo et al., 2020), TransVG (Deng et al., 2021), SimREC (Luo et al., 2022), SeqTR (Zhu et al., 2022), LiteREC (Guan et al., 2023a), RefTrans(Li and Sigal, 2021) and our proposed Potamoi are included. For the RES, we introduce MCN, SeqTR, VLT (Ding et al., 2022), RefTrans and our Potamoi for evaluation. Furthermore, to equally evaluate these models on FVG, we add the same radar backbone and the same fusion methods as Potamoi for image and radar to each model. For our Potamoi, we adopt the BERT tokenizer with a maximum token length of 50. Besides, the pretrained weights of FastViT-SA12 (image-backbone) on ImageNet-1K are loaded during training. The head number of proposed MHSCA are 4, 4 and 8 for three stages while the length of sensor agent $n$ is set as 49.
Training. We resize the input image and RVP map as 640 $\times$ 640 (px). We train all models for 80 epochs with a batch size of 16 based on mixed-precision mode. We set the initial learning rate as 1e-3 with the cosine scheduler. We adopt Adam as the optimizer with a momentum of 0.937 and a weight decay of 5e-4. All training experiments are on four RTX A4000 GPUs.
Evaluation metrics. In contrast to conventional REC and RES tasks where a single text prompt refers to a single target, our WaterVG involves a single text prompt referring to multiple targets. Therefore, for the REC task, we choose to jointly evaluate using $AP_{50}$ (Average Precision at IoU 0.5), $AP_{50-95}$ and $AR_{50-95}$ (Average Precision and Recall from IoU 0.5 to 0.95). $AP_{50}$ considers the localization accuracy above 0.5 IoU confidence, while $AP_{50-95}$ is the average accuracy over a range of IoU thresholds from 0.5 to 0.95. It further considers the localization accuracy under different IoU thresholds. $AR_{50-95}$ is the average recall over the IoU threshold from 0.5 to 0.95. It focuses on the coverage of the model to the true target. Meanwhile, for RES tasks, we adopt mIoU to evaluate models. In tables for comparison, bold text indicates the best, and underlined text indicates the second best.
Table 1. Overall comparison of models on WaterVG.
Vision-based Visual Grounding (VVG)
Methods Val Test
[…]
MattNet 59.6 31.2 39.7 54.81 57.2 30.2 38.9 54.72
MCN 69.9 39.3 50.7 60.16 66.7 39.4 46.2 59.05
TransVG 70.1 40.7 51.0 - 67.2 39.8 48.0 -
SimREC 56.7 31.3 39.8 - 54.0 30.1 37.9 -
SeqTR 71.2 42.0 51.8 62.98 69.1 41.7 49.2 60.03
LiteREC-S2 65.5 37.9 44.9 - 63.5 36.3 46.3 -
RefTrans 70.8 42.8 52.9 63.72 70.8 42.5 50.1 61.03
VLT - - - 64.13 - - - 61.79
Potamoi 71.6 43.3 54.9 63.16 70.0 42.6 50.7 60.36
Fusion-based Visual Grounding (FVG)
MattNet 63.9 38.1 44.0 57.03 59.1 32.3 38.9 60.32
MCN 70.6 45.6 52.5 61.72 68.1 42.5 46.9 66.23
TransVG 71.8 47.1 54.0 - 69.8 43.6 52.0 -
SimREC 60.5 36.8 42.5 - 56.1 30.9 39.7 -
SeqTR 72.6 47.7 54.2 64.96 70.0 43.5 52.6 68.79
LiteREC-S2 68.9 42.7 47.6 - 64.2 36.3 44.5 -
RefTrans 73.1 47.1 53.8 66.19 69.7 43.1 52.7 70.19
VLT - - - 67.59 - - - 70.90
Potamoi 72.8 47.5 55.0 66.27 70.1 44.8 53.0 70.81
5.2. Quantitative Results
Table 1 presents the overall performances of our proposed Potamoi and other competitors on the benchmarks of VVG and FVG. Firstly, overall, we find that all models perform better on FVG than on VVG. This suggests that relying solely on visual features may not accurately localize specific described targets, particularly those incorporating distance or motion features in textual descriptions. Secondly, our proposed Potamoi achieves competitive performance compared to other visual grounding models, whether based on transformer architecture (TransVG, SeqTR, RefTrans) or one-stage (MCN, SimREC, LiteREC-S2) or two-stage paradigms (MattNet). Specifically, for the REC task of VVG and FVG, our Potamoi achieves state-of-the-art performances overall, which in a race with RefTrans (Li and Sigal, 2021) and SeqTR (Zhu et al., 2022). Moreover, for RES task, due to the lightweight segmentation head (Peng et al., 2022) utilized in our Potamoi, it generally ranks second or third among all models in Table 1, trailing behind the dedicated RES model VLT (Ding et al., 2022) based on transformer architecture with a larger parameter count. Nevertheless, from the perspective of multi-task models, our Potamoi overperforms MCN, SeqTR and is evenly contested to RefTrans but a little better.
5.3. Qualitative Results
Figure 8 visualises REC and RES results of our proposed Potamoi and MCN (Luo et al., 2020) on the representative samples of WaterVG, including (a) piecewise multi-objective prompt, (b) distinguish one of multiple objects based on the description of the target distance, (c) multi-object localization based on fuzzy text prompts and (d) count-based fuzzy hints. It can be observed that our proposed Potamoi achieves more prominent performance compared to MCN, obtaining lower target miss rates, false positive rates, and masks that better fit the shapes and contours of the targets.
5.4. Ablation Experiments
Table 2. Text-sensor fusion methods in Potamoi on the test set.
[…]
Stage 2: $D$=128, $N_{h}$=4, $H$=80, $W$=80, $L$=50
Stage 3: $D$=256, $N_{h}$=4, $H$=40, $W$=40, $L$=50
Stage 4: $D$=512, $N_{h}$=8, $H$=20, $W$=20, $L$=50
MHCA (Dao et al., 2022) 66.048K 293.27M 88.00 71.8 70.23
0.263M 257.23M 64.06
1.050M 256.41M 60.00
MHLCA (Choromanski et al., 2020) 0.252M 1.691G 124.02 70.2 72.01
0.524M 865.08M 94.27
1.049M 471.86M 45.74
Default Length of Sensor Agent $n$ = 49 (7 $\times$ 7)
MHSCA 0.768K 8.11M 81.48 70.1 70.81
1.536K 3.98M 49.83
3.072K 1.99M 34.77
Length of Sensor Agent $n$ = 144 (12 $\times$ 12)
MHSCA 0.768K 8.14M 108.53 71.2 71.04
1.536K 4.07M 66.14
3.072K 2.04M 44.29
Cross attention comparison. In Table 2, we conduct a comprehensive comparison between our MHSCA and two representative multi-head cross-attention mechanisms, MHCA (Dao et al., 2022) and MHLCA (Choromanski et al., 2020), which are transplanted to our Potamoi, encompassing comparisons across different channel dimensions and spatial sizes of visual feature maps. Moreover, we adjust the agent number $n$ in MHSCA as 49 and 144 for comparison, respectively.
Firstly, looking at the performance on both REC and RES tasks, our proposed MHSCA ($n$=144) achieves competitive results. It outperforms MHLCA slightly in the REC task and lags behind MHCA, while in the RES task, it surpasses MHCA but still trails MHLCA. However, it is noteworthy that for visual feature maps with increasing spatial sizes, the FLOPs of our MHSCA never exceed 10M and demonstrate smaller growth compared to MHCA and MHLCA. Regarding expanding channels, our MHSCA maintains a significantly lower parameter count compared to the other two fusion methods. Importantly, we also compare the power consumption of three fusion methods, where our MHSCA with $n$=49 achieves the lowest value for three-dimensional feature maps. MHSCA with $n$=144 obtains competitive performance with MHCA based on flash attention (Dao et al., 2022). Therefore, our MHSCA achieves a well-balanced multi-modal fusion approach between accuracy, complexity and power consumption, serving as a lightweight cross-attention mechanism.
Comparison of other settings. As Table 3 shows, when adopting multi-task learning, the performance of Potamoi is apparently better than single-task learning, which indicates multi-task learning can jointly promote the performances of individual tasks. Figure 9 visualizes the last stage of P4 in the Lite-PAN before REC and RES heads, we find that Potamoi successfully focuses on the spatial locations of one or more targets referred to in textual prompts, indicating that REC and RES are optimized in the right direction. Table 4 presents the ablation experiments of other settings in fusion-based Potamoi, mainly including the settings of input data (RVP maps) and neural networks. We find that in the radar features of targets, target distance and velocity are relatively more vital for visual grounding. Additionally, due to some samples having text prompt token lengths exceeding 30, and certain target-descriptive terms occurring towards the end of text, reducing token length resulted in information loss, leading to a decrease in prediction accuracy. Moreover, Adaptive Radar Weighting (ARW), as well as SPPM, play significant roles on feature representation.
Table 3. Ablation experiments on multi-task learning.
[…]
70.9 44.0 52.6 70.68
70.1 44.8 53.0 70.81
Table 4. Ablation experiments on other settings of Potamoi.
[…]
Baseline 70.1 44.8 53.0 70.81
Input Data
Radar map (R) 70.0 42.8 52.5 70.41
Radar map (R-V) 70.1 43.1 52.4 70.55
Radar map (R-P) 69.8 42.9 52.4 70.24
Radar map (V-P) 68.6 42.7 51.9 68.23
Token Len=15 67.8 41.5 45.9 65.79
Token Len=30 69.8 43.2 52.0 70.31
Network Settings
No ARW 69.5 42.8 52.2 70.40
No SPPM - - - 69.52
No Query Pos 68.9 42.7 52.0 69.97
Figure 9. Last-stage heatmaps of PAN before prediction heads.
6. Conclusion
In this paper, we propose a new visual grounding dataset, WaterVG, for intelligent perception in waterway environments, which includes textual prompts, visual, and millimeter-wave radar data. Additionally, each sample’s text prompt contains references to multiple targets. Based on this, we design an efficient multitask visual grounding model, Potamoi. It is capable of accurately performing REC and RES tasks simultaneously based on textual prompts containing target appearance, position, distance features, and more. This capability relies on our proposed lightweight and highly effective Multi-Head Slim Cross Attention (MHSCA) for effective fusion of textual, visual, and radar features. Both Potamoi and MHSCA demonstrate competitive performance compared to counterparts.