### v2

# WaterVG: Waterway Visual Grounding based on Text-Guided Vision and mmWave Radar

Abstract.
The perception of waterways based on human intent is significant for autonomous navigation and operations of Unmanned Surface Vehicles (USVs) in water environments. Inspired by visual grounding, we introduce WaterVG, the first visual grounding dataset designed for USV-based waterway perception based on human prompts. WaterVG encompasses prompts describing multiple targets, with annotations at the instance level including bounding boxes and masks. Notably, WaterVG includes 11,568 samples with 34,987 referred targets, whose prompts integrates both visual and radar characteristics. The pattern of text-guided two sensors equips a finer granularity of text prompts with visual and radar features of referred targets. Moreover, we propose a low-power visual grounding model, Potamoi, which is a multi-task model with a well-designed Phased Heterogeneous Modality Fusion (PHMF) mode, including Adaptive Radar Weighting (ARW) and Multi-Head Slim Cross Attention (MHSCA). Exactly, ARW extracts required radar features to fuse with vision for prompt alignment. MHSCA is an efficient fusion module with a remarkably small parameter count and FLOPs, elegantly fusing scenario context captured by two sensors with linguistic features, which performs expressively on visual grounding tasks. Comprehensive experiments and evaluations have been conducted on WaterVG, where our Potamoi archives state-of-the-art performances compared with counterparts.
Waterway visual grounding, Multi-modal fusion, Low-power model, USV-based perception, mmWave radar
submissionid: 687
Figure 1. Overview of our pipeline for WaterVG. The camera and radar align prompt on appearance, motion and distance features.
1. Introduction
With the advancement of deep learning, water-surface intelligent perception has witnessed substantial progress in various domains (Yao et al., 2023a, b, c; Guo et al., 2023). Specifically, current advancements primarily encompass object detection (Cheng et al., 2021c; Guan et al., 2023d), segmentation(Guan et al., 2023c, b, 2024), tracking (Zhang et al., 2021) and Simultaneous Localization And Mapping (SLAM) (Cheng et al., 2021a). Based on the aforementioned, waterway perception faces the following challenges, firstly, waterways are often characterized by complexity, irregularity, and variability in obstacles and lighting conditions. Secondly, due to fluctuations in water levels and changes in flow velocity, many obstacles are irregular and shifting. Thirdly, vessel navigation in waterways usually does not follow prescribed routes, leading to a higher frequency of unforeseen events and increased challenges in waterway monitoring. Therefore, it is challenging for regulators or specialized USVs to locate desired waterway targets.
However, current sensor-based full-target perception may not adequately address this issue. Hence, we pioneeringly introduce visual grounding to waterway perception, utilizing natural language as prompt to locate specific targets. Here, humans can convey target exact features through descriptions, facilitating more intuitive and efficient guidance for USVs. This leads to enhanced precision in navigation and localization, thereby holding significant implications for transportation, rescue operations, and environment conservation.
Figure 2. Samples in WaterVG. The annotations include bounding boxes and masks, supplemented by radar point clouds (red).
Building upon the aforementioned considerations, we contemplate the establishment of an optimal visual grounding dataset tailored for robust perception and monitoring in waterway environments. Here are our considerations, firstly, pure visual sensors are easily disturbed by variable visibility and lighting conditions (Cheng et al., 2021c; Yao et al., 2023a, c). Secondly, visual sensors can only capture qualitative features of targets in terms of appearance or position, but traffic supervisors often need quantitative distance and motion information about waterway targets. Therefore, when one desires to ascertain whether there is a vessel approaching behind the fog or whether obstacles exist within a specific distance, visual sensors are unable to precisely locate targets based on human intent. To enable target referring flexibly, we complement the visual sensor with 4D mmWave radar, as it is an all-weather sensor that can detect targets in the form of point clouds and offer higher resolution than conventional mmWave radar (Guan et al., 2023b, d; Cheng et al., 2021b). Concurrently, while visual sensors perceive appearance features, including color, size, texture and shape, radar can furnish target reflection characteristics (Yao et al., 2023b), including distance, velocity, reflected power, etc.
Upon these insights, we build the first visual grounding dataset called WaterVG for waterway perception, based on the WaterScenes dataset (Yao et al., 2023c), which is a USV-based waterway perception dataset integrating the monocular camera and 4D radar. Detailedly, diverging from other vision-based visual grounding datasets, our WaterVG dataset has the following distinguishing features: i) Manual fine-grained prompts to describe targets through both visual attributes like color, size, shape, texture and features detectable by radar, such as numeric distance and motion characteristics. ii) Flexible referred targets with 1 to 15 per prompt. iii) Diverse prompts, containing specific descriptions, coarse and fine queries on single or multiple characteristics. iv) Multi-task annotations of Referring Expression Comprehension (Qiao et al., 2020) and Segmentation (Chrupała, 2022) (REC and RES).
Correspondingly, three heterogeneous modalities, RGB image, 3D point clouds and natural language will face several challenges regarding fusion for USV-based visual grounding. Firstly, radar may be interfered with clutter, causing some non-target points (ghost echoes). Hence, it is crucial to adaptively extract radar features mentioned in prompts while ignoring clutter. Secondly, referred targets vary significantly in visual and point cloud scale, so the dynamic fusion of sparse sensor features with corresponding dense language features is vital. Thirdly, fragmented feature extraction and fusion branches of multi-modal networks tend to cause increased power consumption and reduction in USV endurance time.
Therefore, we propose a low-power multi-task model called Potamoi, achieving a trade-off between accuracy and power consumption. Within this, we design an effective fusion mode called Phased Heterogeneous Modality Fusion (PHMF). PHMF performs a phased alignment and fusion of visual, radar, and linguistic features. In PHMF, Adaptive Radar Weighting (ARW) is proposed for high-quality radar feature extraction, effectively suppressing interference from clutter in space and adaptively extracting channel features of target radar characteristics contained within text prompts to the maximum extent. More importantly, PHMF’s core lies in our proposed Multi-Head Slim Cross Attention (MHSCA). MHSCA efficiently reduces redundancy in query context and fuses contextual features captured by two sensors with linguistic features, achieving exemplary performance on visual grounding with lower parameters and FLOPs than vanilla Multi-Head Cross Attention (MHCA) (Wu et al., 2023) and Multi-Head Linear Cross Attention (MHLCA) (Choromanski et al., 2020).
In summary, the contributions of this paper are as follows:
  1. (1)
    The first visual grounding dataset for USV-based waterway perception, WaterVG, comprises 11,568 triplet input data, including images, radar point clouds, and prompts. Moreover, the ground truth encompasses two types of instance-level annotations: bounding boxes and masks. In particular, our textual prompts contain extra corresponding descriptions of target distance and motion features.
  2. (2)
    A low-power multi-task visual grounding model, Potamoi, where an effective fusion mode called PHMF, aligns and fuses three modalities. PHMF includes ARW for radar feature adaptive extraction and a low-cost fusion method called MHSCA, which can fuse heterogeneous modalities efficiently for visual grounding with few parameters, low FLOPs and power than vanilla MHCA and MHLCA.
  3. (3)
    Comprehensive experiments based on the WaterVG dataset encompass various aspects such as the performance of visual grounding models under fusion and vision patterns, and multi-modal fusion methods. These experiments deepen insights and understanding of the field.
Figure 3. The statistics of WaterVG, including proportion on the proportion of waterways (a), sentence patterns (b) and query types (c), and the statistics of WaterVG on referred target number (d), category distribution (e), prompt length (f) and word cloud (g).
2. Related Works
2.1. Waterway Perception Datasets
Recently, waterway perception datasets have obtained significant attention in the field of waterway robotics and autonomous navigation. Most waterway perception datasets cover instance-level water-surface obstacle detection (Kristan et al., 2015; Bovcon et al., 2018; Liu et al., 2021; Cheng et al., 2021c), segmentation (Bovcon et al., 2019; Taipalmaa et al., 2019; Zhou et al., 2021; Žust et al., 2023) and tracking (Moosbauer et al., 2019; Varga et al., 2022) for traffic monitoring, autonomous driving and water environmental protection. Furthermore, datasets for the SLAM (Miller et al., 2018; Cheng et al., 2021a) and segmentation (Yao et al., 2023c; Cheng et al., 2021a) of drivable-area and shorelines are also proposed for the planning-oriented perception of USVs. Based on the above observations, it is evident that current waterway perception datasets primarily focus on sensor-based automatic perception, lacking flexibility in human-machine interaction. Text prompts can be regarded as a way to simulate human instructions. In many scenarios, leveraging text prompts, USVs (waterborne robots) can more flexibly locate and identify specific targets, thereby achieving more efficient execution of waterway tasks. Therefore, we design and construct the first waterway visual grounding dataset, WaterVG, to promote human-guided waterway monitoring and USV-based embodied AI in waterway perception.
2.2. Visual Grounding Datasets and Models
As a challenging branch within multi-modal learning, visual grounding aims to locate, based on text query, the targets in images that best match the descriptions provided by the natural language (Zhou et al., 2023). This is further divided into REC and RES tasks. The REC task focuses on localizing targets with bounding boxes, whereas RES aims to delineate target masks. The majority of current 2D visual grounding datasets (Kazemzadeh et al., 2014; Plummer et al., 2015; Yu et al., 2016; Krishna et al., 2017; Deruyttere et al., 2019; Wang et al., 2020; Chen et al., 2020; Zhan et al., 2023; Liu et al., 2023; Wu et al., 2023; He et al., 2023) primarily cater to REC tasks, relying solely on pure visual data, lacking features from other synchronous modalities queried by prompts. Moreover, there is a current trend towards a prompt capable of querying multiple targets, which aligns more flexibly with real-world scenarios, deviating from the single-target reference. Moreover, visual grounding models are primarily categorized into three classes: two-stage (Liu et al., 2019; Yang et al., 2019b; Yu et al., 2018; Wang et al., 2019; Hong et al., 2019), one-stage (Luo et al., 2020; Chen et al., 2018; Yang et al., 2019a), and transformer-based models (Zhu et al., 2022; Du et al., 2022; Deng et al., 2021). Among them, firstly, two-stage models split the generation of Regions Of Interest (ROI) and target classification into two parts, where the first stage is responsible for matching the regions of interest with text features while the second stage is to filter and refine ROI based on the fused features of text and image. Secondly, one-stage models usually combine bounding box regression and confidence prediction into one step. Besides, linguistic features usually fuse with image features at the end of the image backbone or within multi-scale fusion networks. Thirdly, transformer-based models are usually in the format of an image backbone and a text encoder followed by a transformer-based encoder and decoder, where the fusion operation usually happens before the transformer-based encoder.
3. The WaterVG Dataset
Overall, the WaterVG dataset comprises 11,568 samples, including 9,254 in the training set, 578 in the validation set, and 1,736 in the test set. Each sample consists of an image, a frame of 4D radar point cloud, a prompt description, and ground truth presented in the forms of bounding boxes and masks. This section will sequentially elaborate on the construction process (Section 3.1) and data statistics (Section 3.2) of the WaterVG dataset.
Figure 4. Five-step prompt annotations of WaterVG.
Figure 5. The architecture of Potamoi. The text encoder (ALBERT) (Lan et al., 2020) for text prompts is frozen during training.
3.1. Dataset Construction
Our WaterVG dataset is built upon the WaterScenes dataset (Yao et al., 2023c), a high-quality, multi-sensor dataset with vision and radar, specifically designed for various water-surface environments. It features high-resolution images from a monocular camera with 1920 × 1080 pixels, complemented by 4D radar point clouds, which are calibrated and time-synchronized with the monocular camera. Additionally, WaterScenes (Yao et al., 2023c) provides high-quality annotations for object detection in the form of bounding boxes, as well as segmentation masks. Building upon these resources, we engage two experts with extensive experience in intelligent waterway transportation to conduct prompt annotation of our WaterVG dataset, selecting 11,568 representative samples from WaterScenes under different scenarios and lighting conditions. As illustrated in Fig. 4, the annotation process is divided into five distinct stages.
(1) Select candidate targets. To facilitate efficient annotation by experts, we develop an annotation GUI client specifically for the WaterVG dataset. In this system, the radar point clouds are projected onto the camera plane using the projection matrix based on the extrinsic parameters of the radar and camera, as well as the intrinsic parameters of the camera, with the point clouds displayed in red. Each point cloud is accompanied by information displaying the distance of the target from the USV and its radial velocity. Correspondingly, both bounding box and mask annotations are mapped onto the image. Annotators simply need to click on the location of the target to display these two types of annotations.
(2) Check the annotations. Once the bounding box and mask of a clicked object are displayed on the image, two annotators proceed to check the accuracy of corresponding ground truth. If the ground truth is verified as correct, they continue the next step. If not, the target is re-annotated to ensure the precision of the dataset.
(3) Write referring prompts. Based on the images and radar point clouds, two annotators include as many physical characteristics of the targets as possible in the annotation prompt. These encompass the target’s category, size, color, and shape features perceptible by visual means. Furthermore, if the target is detected by the radar precisely accompanied by dense projected point clouds, we use characteristics of point clouds to further describe the target’s distance and motion trends from our USV. Additionally, if multiple targets are involved, annotators will also work on expressing their spatial location relationships. Based on these, the annotators aim to diversify the prompts by incorporating various sentence structures and methods of referencing the target.
(4) Enrich prompts. When facing identical scenes or targets in different samples, the limited vocabulary and subjective tone of annotators are unavoidable. To enhance the diversity of prompt content, we utilize the existing annotated prompts and feed them into ChatGPT-3.5-Turbo API (Ouyang et al., 2022), aiming to enrich the prompts by generating semantically equivalent yet more varied expressions.
(5) Check each other’s prompt annotations. To avoid ambiguity in prompts, we exchange prompts written by two experts. Each expert reviews and examines prompts written by the other, respecting respective writing styles. In cases where there is ambiguity or misidentification in target references, the two experts engage in discussions and reach a consensus for necessary modifications.
3.2. Dataset Statistics
Figure 3 presents the statistics of WaterVG from perspectives of waterway, sentence pattern and query type. The waterway scenario proportion in WaterVG, includes canal, river, lake and moat. Besides, sentence patterns for prompts, including statement (egocentric), question, statement (imperative) and exclamation. Prompt type proportion contains 4 main types of query by exact features, partial features, number and category, respectively. Figure 3 also illustrates the distribution statistics of referred target number, prompt length per sample, referred category distribution and word cloud. In WaterVG, firstly, each text prompt refers to 1-15 targets, with the majority consisting of 1-2 targets. Secondly, The categories of referred target include piers, buoys, sailors, ships, boats, pleasure boats, kayaks and rubbishes, a total of 34,987 referred targets. Thirdly, the distribution of text prompt lengths, ranging from 2 to 40, encompass simple phrases and detailed sentences. Fourthly, the word cloud displays high-frequency words. The above demonstrates the diversity and comprehensiveness of our WaterVG dataset regarding waterway visual grounding, which is primarily reflected in the diversity of waterway scenes, richness of textual prompts in terms of content, patterns, and types, diversity of mentioned target categories, and flexibility in mentioned target number.
4. Methods
4.1. Task Formulation
Potamoi is a one-stage multi-task visual grounding model for two tasks, REC and RES. Given triple inputs of an image $I\in R^{3\times H_{I}\times H_{W}}$, a frame of radar point clouds $RPC\in R^{5\times N_{R}}$, containing the features $\{x,y,z,v,p\}$ where $v$ and $p$ denote compensated velocity and reflected power, and a text prompt sequence $T=\{w_{1},w_{2},\dotsc,w_{i}\}$ with a length of $N_{T}$. The output of REC is a set of bounding boxes. The $i$-th bounding box is formulated as $B_{i}=\{x_{i},y_{i},w_{i},h_{i},c_{i}\}$, which represents the center coordinates, width, height and IoU confidence of bounding box. RES is to predict a mask of referred targets $M_{RES}\in R^{2\times H_{I}\times H_{w}}$, where the channel dimension contains referred foreground target and unreferred background.
4.2. Overall Pipeline
Feature Encoders. Firstly, the image encoder adopts FastViT-SA24 (Vasu et al., 2023) with the cascade of reparametrized RepMixer and Multi-Head Self-Attention. Secondly, to prepare for fusion with image, we project 3D radar point clouds from radar to image coordinate, and obtain a 3-channel RVP map, where each position contains the target Range (R), compensated Velocity (V) and reflected Power (P). We adopt the combination of convolution, batch normalization and ReLU activation as the basic block to exact radar features. Each stage of radar encoder contains three blocks, where the convolution kernels are $3\times 3$, $1\times 1$ and $3\times 3$ with the stride of $1$, $1$ and $2$ for local feature extraction, feedforward and downsampling. Whatever image or radar encoder, there are one stem layer and four stages, and the stem layer downsamples the spatial size four times. The feature dimension of stage $i\in\{1,2,3,4\}$ is $(64\times 2^{i-1})\times\frac{I_{H}}{2^{i+1}}\times\frac{I_{W}}{2^{i+1}}$.
Feature Fusion. To ensure precise guidance of prompt for both sensors to locate referred targets, the fusion with text requires high-quality integration of features from both sensors. Hence, we adopt a divide-and-conquer strategy called Phased Heterogeneous Modality Fusion. Firstly, due to water-surface reflection and multi-path effect from buildings on the riverbank, radar features unavoidably contain clutter in some scenarios, which results in a certain level of interference in fused sensor features. Secondly, not all prompts contain radar features of targets, so adaptive channel weighting matters. To tackle above problems, a lightweight module called Adaptive Radar Weighting is devised for clutter alleviation, where the last three stage features from the radar encoder will first go through three ARW modules before fusing with image features. Then, three-stage sensor features act as queries to fuse with the keys and values of textual features from the pretrained ALBERT (Lan et al., 2020) by our proposed low-power MHSCA and output three fusion features, which mitigates the quadratic complexity and high power consumption caused by vanilla MHCA (Wu et al., 2023). A lite Path Aggregation Network (PANet) (Li et al., 2022) is to fuse features with various receptive fields, and outputs four-stage pyramid features from $P2$ to $P5$.
Prediction Heads. We feed all four pyramid features to REC heads, which are adapted upon YOLOv8 decouple head (Jocher et al., 2023). Each pyramid feature corresponds to one REC head, containing confidence and bounding box regression branch. The shape of REC output for the stage $i\in\{2,3,4,5\}$ is $(1+4\times RegMax)\times\frac{I_{H}}{2^{i+1}}\times\frac{I_{W}}{2^{i+1}}$, where $1$ is the confidence score and $4$ is four coordinates of bounding box. $RegMax$ is the hyperparameter of Distributed Focal Loss (DFL) (Li et al., 2020). Further, following the paradigm of PP-LiteSeg (Peng et al., 2022), we feed $P2$, $P3$ and $P4$ with SPPM (Jocher et al., 2023) to the RES head, which is based on FLD (Peng et al., 2022). The output shape of RES head is $2\times I_{H}\times I_{W}$, including referred foreground and unreferred background categories.
4.3. Phased Heterogeneous Modality Fusion
Figure 6. The structure of Adaptive Radar Weighting.
Figure 7. The structure of Slim Cross Attention.
(a) Adaptive Radar Weighting. ARW is employed to dynamically weigh radar features on both channels and space, obtaining high-quality sensor features fused upon vision and radar features while adaptively aligning useful radar features within characteristics of targets contained in the text prompts.
As Figure 6 presents, assuming the radar feature is $F_{R}\in R^{C\times H\times W}$, we exert a linear projection with GeLU activation on it to obtain $\hat{F_{R}}\in R^{C\times H\times W}$. The resulting feature map is then prepared for adaptive channel attention matrix $W_{CA}\in R^{C\times 1\times 1}$ by applying adaptive average pooling to collapse spatial dimensions and computing attention weights by softmax. $\hat{F_{R}}$ is position-wise multiplication with $W_{CA}$ to obtain the channel-weighting feature $F_{R}^{CW}\in R^{C\times H\times W}$. Then spatial weighting matrix $W_{SA}\in R^{1\times H\times W}$ is achieved by convolving $F_{R}^{CW}$ with a GELU activation and a 1 $\times$ 1 convolution, followed by element-wise multiplication with the original features and obtain the final radar feature $F_{R}^{ARW}\in R^{C\times H\times W}$ with both channel and spatial adaptive weighting. Equation 1 and 2 present the whole process above.
[…]
(b) Multi-Head Slim Cross Attention. After obtaining three-stage sensor features ($F^{2}_{S}$, $F^{3}_{S}$ and $F^{4}_{S}$) containing contextual features of vision and prior target features captured by radar, there are three linear projection branches from ALBERT to align the channel dimensions of three-stage sensor features. To effectively fuse two heterogeneous modalities with a low cost on USV-based edge devices, we propose Multi-Head Slim Cross Attention (MHSCA) (Figure 7). Given a sensor feature $F_{S}\in R^{C\times H\times W}$ and a text feature $F_{T}\in R^{L\times C}$, we first exert learnable encodings (Dosovitskiy et al., 2020)(Vaswani et al., 2017), $P_{I}$ and $P_{T}$ on them, respectively. Then we adopt 2D depth-wise convolution with $1\times 1$ kernel ($W_{Q}$) to obtain the sensor query ($Q$) while two 1D depth-wise convolutions ($W_{K}$ and $W_{V}$) are for generate key ($K$) and value ($V$). Equation 3 presents the aforementioned process.
(3) $Q=F_{S}W_{Q}+P_{S},K=F_{T}W_{K}+P_{T},V=F_{V}W_{V}+P_{T},$
$Q\in R^{C\times H\times W},K\in R^{L\times C},V\in R^{L\times C},$
Further, we employ adaptive pooling to downsample the sensor query $Q$ (Han et al., 2023), preserving its essential context while reducing its spatial size to obtain the sensor agent $A\in R^{C\times h\times w}$. As the agent of $Q$, $A$ is firstly flattened along spatial dimension as $\hat{A}\in R^{n\times C}$ with $n=h\times w$, and conducts softmax attention $\sigma(\cdot)$ with $K$ and $V$ to calculate the agent-conditional Value matrix $V_{A}$. (Equation 4, 5)
(4) $\hat{A}=Flat(A)^{T},\hat{A}\in R^{n\times C},$
(5) $V_{A}=\sigma(\hat{A},K,V)=\frac{\hat{A}K^{T}}{\sqrt{d}}\cdot V,$
where $d$ is the channel dimension of $F_{S}$ and $F_{T}$.
Then, $V_{A}$ is forwarded to the second softmax attention, broadcasting the inter-modal similarity embedded in text feature to the sensor query $Q$, where $A$ acts as the key and $V_{A}$ is the value (Equation 6, 7). Here $Q$ is also firstly flattened as $\hat{Q}\in R^{N\times C}$ with $N=H\times W$.
(6) $\hat{Q}=Flat(Q)^{T},\hat{Q}\in R^{N\times C},$
(7) $F_{TS}=\sigma(\hat{Q},A,V_{A})=\frac{\hat{Q}A^{T}}{\sqrt{d}}\cdot V_{A},$
After that, a residual path from the sensor feature $F_{S}$ is added to the output of the second softmax attention and obtains the image-like sensor-conditional text features $F_{TS}$ (Equation 8). Finally, a 2D depth-wise convolution with $1\times 1$ kernel size ($W$) is exerted on $F_{TS}$ to further feedforward.
(8) $\hat{F}_{TS}=(Reshape(F_{TS})+F_{S})W,F_{TS}\in R^{C\times H\times W},$
Here, the computation complexity of our proposed slim cross attention is $O(Nnd)+O(NLd)$, which is lower than vanilla cross attention’s (Wu et al., 2023) $O(NLd)+O(NLd)$, as $n\ll L$ in practice.
4.4. Training Objectives
We reformulate the REC and RES tasks as the bounding box regression and semantic segmentation. Though all targets are denoted as one category, referred target, implicit unbalance between target categories still exist. Hence, as Equation 9 presents, we introduce Distributed Focal Loss (DFL) (Li et al., 2020) to alleviate the unbalanced distribution while CIoU (Zheng et al., 2020) is added to refinedly regress the bounding box size by taking its aspect ratio into account.
(9) $L_{REC}=L_{DFL}+L_{CIoU},$
For RES task, we combine the dice loss (Sudre et al., 2017) with focal loss (Lin et al., 2017) to jointly subtly optimize the mask prediction (Equation 10).
(10) $L_{RES}=L_{Dice}+L_{FL},$
We empirically argue that the loss functions are not in the same order of magnitude for REC and RES during multi-task training. To avoid the domination of optimization by the task with a large loss, we introduce uncertainty weighting (Kendall et al., 2018) to balance multi-task optimization. Equation 11 demonstrates the loss function during the joint optimization of REC and RES.
(11) $L_{Joint}=\frac{1}{2\sigma_{1}^{2}}L_{REC}+\frac{1}{2\sigma_{2}^{2}}L_{RES}+log\sigma_{1}+log\sigma_{2},$
where $\sigma_{1}$ and $\sigma_{2}$ are learnable parameters to weigh two loss items.
5. Experiments
5.1. Settings of Dataset and Implementation
Datasets. Our proposed WaterVG is for training, validation and testing on all models. WaterVG includes two benchmarks, Fusion-based Visual Grounding (FVG) and Vision-based Visual Grounding (VVG). Moreover, to comprehensively valid the performance of Potamoi, we also evaluate our Potamoi on RefCOCO (testA, testB), RefCOCO+ (testA, testB) and RefCOCOg (test-umd) (Yu et al., 2016).
Models. For the REC, MattNet (Yu et al., 2018), MCN (Luo et al., 2020), TransVG (Deng et al., 2021), SimREC (Luo et al., 2022), SeqTR (Zhu et al., 2022), LiteREC (Guan et al., 2023a), RefTrans(Li and Sigal, 2021) and our proposed Potamoi are included. For the RES, we introduce MCN, SeqTR, VLT (Ding et al., 2022), RefTrans and our Potamoi for evaluation. Furthermore, to equally evaluate these models on FVG, we add the same radar encoder and ARW as Potamoi for image and radar to each model. For Potamoi, we adopt ALBERT tokenizer (Lan et al., 2020) with a maximum token length of 50. Besides, the pretrained weights of FastViT-SA24 (image encoder) and other models’ backbones on ImageNet-1K (Deng et al., 2009) are loaded during training. The head number of proposed MHSCA are 4, 4 and 8 for three stages while the length of sensor agent $n$ is set as 49. On RefCOCOs, our Potamoi is trained from scratch.
Training. We resize the input image and RVP map as 640 $\times$ 640 (px). We train all models for 80 epochs on WaterVG with a batch size of 16 while 100 epochs with 64 batch size on RefCOCOs by mixed-precision mode. We set the initial learning rate as 1e-3 with the cosine scheduler. We adopt SGDM as the optimizer with a momentum of 0.937 and a weight decay of 5e-4. All training experiments are on four RTX A4000 GPUs.
Evaluation metrics. Compared with datasets (Yu et al., 2016; Plummer et al., 2015) where a prompt refers to one target, our WaterVG involves a single text prompt referring to multi-targets. Hence, to comprehensively evaluate model performances, we use $AP_{50}$ (Average Precision above IoU 0.5), $AP_{50-95}$ and $AR_{50-95}$ (Average Precision and Recall from IoU 0.5 to 0.95) in REC. For RES tasks, we adopt mIoU to evaluate models. Moreover, we evaluate models on their power consumption (Tu et al., 2023). In tables for comparison, bold text indicates the best, and underlined text indicates the second best. For RefCOCOs, we use $AP_{50}$ as the metric of REC while mIoU as the RES metric.
Figure 8. Images with projected radar point clouds (PCs) (first row) and prediction results by Potamoi (second row) and MCN (Luo et al., 2020) (third row) under different situations. Red bounding boxes are REC predictions while translucent red masks are RES predictions. Blue PCs denote clutter while others represent targets. The gray letters on PCs refers to the distance (m) from targets to the ego USV.
Table 1. Overall comparison of models on WaterVG. MT means multi-task while $P$ is power consumption calculated by hardware API.
Vision-based Visual Grounding (VVG) Fusion-based Visual Grounding (FVG)
Methods MT Val Test Val Test
[…]
MattNet 59.6 31.2 39.7 54.81 57.2 30.2 38.9 54.72 - 63.9 38.1 44.0 57.03 59.1 32.3 38.9 60.32 -
MCN 69.9 39.3 50.7 60.16 66.7 39.4 46.2 59.05 81.34 70.6 45.6 52.5 61.72 68.1 42.5 46.9 66.23 77.72
TransVG 70.1 40.7 51.0 - 67.2 39.8 48.0 - - 71.8 47.1 54.0 - 69.8 43.6 52.0 - -
SimREC 56.7 31.3 39.8 - 54.0 30.1 37.9 - 94.42 60.5 36.8 42.5 - 56.1 30.9 39.7 - 88.70
SeqTR 71.2 42.0 51.8 62.98 69.1 41.7 49.2 60.03 90.61 72.6 47.7 54.2 64.96 70.0 43.5 52.6 68.79 84.65
LiteREC 65.5 37.9 44.9 - 63.5 36.3 46.3 - 70.22 68.9 42.7 47.6 - 64.2 36.3 44.5 - 65.79
RefTrans 70.8 42.8 52.9 63.72 70.8 42.5 50.1 61.03 92.97 73.1 47.1 53.8 66.19 69.7 43.1 52.7 70.19 87.21
VLT - - - 64.13 - - - 61.79 - - - - 67.59 - - - 70.90 -
Potamoi 71.6 43.3 54.9 63.16 70.0 42.6 50.7 60.36 66.10 72.8 47.5 55.0 66.27 70.1 44.8 53.0 70.81 61.48
Table 2. Potamoi under waterways, sentence and query patterns.
Waterways Canal Moat River Lake
AP50 72.5 69.4 70.5 67.5
mIoU 72.90 71.81 73.45 66.48
Sentences SE1 SI2 QU3 EX4
AP50 73.6 70.0 65.5 70.1
mIoU 74.12 71.53 66.41 68.74
Queries Exact-Feat Partial-Feat Number Category
AP50 73.1 70.1 62.2 68.7
mIoU 74.35 71.76 63.57 66.87
1. Statement (egocentric), 2. Statement (imperative), 3. Question, 4. Exclamation.
5.2. Quantitative Results
Table 1 presents the overall performances of our proposed Potamoi and other competitors on the benchmarks of VVG and FVG. Firstly, overall, we find that all models perform better on FVG than on VVG. This suggests that relying solely on visual features may not accurately localize specific described targets, particularly those incorporating distance or motion features in textual descriptions. Secondly, our proposed Potamoi achieves competitive performance compared to other visual grounding models, whether based on transformer architecture (TransVG, SeqTR, RefTrans) or one-stage (MCN, SimREC, LiteREC-S2) or two-stage paradigms (MattNet). Specifically, for REC task of VVG and FVG, our Potamoi achieves state-of-the-art performances overall, which in a race with RefTrans (Li and Sigal, 2021) and SeqTR (Zhu et al., 2022). Moreover, for RES task, due to the lite segmentation head (Peng et al., 2022) in Potamoi for power saving, it generally ranks second or third among all in Table 1, trailing behind the dedicated RES model VLT (Ding et al., 2022) based on transformer architecture with a larger parameter count. Nevertheless, among multi-task models, our Potamoi overperforms MCN, SeqTR and is evenly contested to RefTrans. Table 2 presents the performance of Potamoi on test set under different scenarios and sentence query patterns as Figure 3 presents. Importantly, our Potamoi achieves the lowest power consumption for both VVG or FVG, which is much lower than both single-task and multi-task models, implying Potamoi can help extend the endurance time of USV operations. Lastly, Table 4 and 5 show that our low-power Potamoi obtains competitive performances compared with other state-of-the-art models on RefCOCOs. Further, our MHSCA can be nicely adapted to CNN-based MCN and transformer-based SeqTR to improve performance.
Table 3. Text-sensor fusion methods in Potamoi on the test set.
[…]
Stage 2: $D$=128, $N_{h}$=4, $H$=80, $W$=80, $L$=50
Stage 3: $D$=256, $N_{h}$=4, $H$=40, $W$=40, $L$=50
Stage 4: $D$=512, $N_{h}$=8, $H$=20, $W$=20, $L$=50
MHCA (Dao et al., 2022) 66.048K 293.27M 71.8 45.0 52.9 70.23
0.263M 257.23M
1.050M 256.41M
MHLCA (Choromanski et al., 2020) 0.252M 1.691G 70.2 44.7 53.2 72.01
0.524M 865.08M
1.049M 471.86M
Default Length of Sensor Agent $n$ = 49 (7 $\times$ 7)
MHSCA 0.768K 8.11M 70.1 44.8 53.0 70.81
1.536K 3.98M
3.072K 1.99M
Length of Sensor Agent $n$ = 144 (12 $\times$ 12)
MHSCA 0.768K 8.14M 71.7 45.1 53.1 71.09
1.536K 4.07M
3.072K 2.04M
Length of Sensor Agent $n$ = 256 (16 $\times$ 16)
MHSCA 0.768K 8.24M 72.3 45.2 54.0 71.95
1.536K 3.98M
3.072K 2.26M
Figure 9. Prediction samples of Potamoi on four primary query types of textual prompts.
5.3. Qualitative Results
Figure 8 visualizes predictions of our Potamoi and MCN (Luo et al., 2020) on the representative samples of WaterVG, including (a) distinguish partial targets based on the query of target distance, (b) multi-target localization based on fuzzy text prompts and (c) count-based fuzzy hints. Apparently, our Potamoi behaves more prominent than MCN, obtaining lower target miss rates, false positive rates, and masks that better fit shapes and contours of the targets. Moreover, Figure 9 presents prediction samples upon four primary query types in WaterVG, where our Potamoi exhibits nice adaptation.
5.4. Ablation Experiments
Cross attention comparison. In Table 3, we conduct a comprehensive comparison between our MHSCA and two representative multi-head cross-attention mechanisms, MHCA (Dao et al., 2022) and MHLCA (Choromanski et al., 2020), which are transplanted to our Potamoi, encompassing comparisons across different channel dimensions and spatial sizes of visual feature maps in PAN. Moreover, we adjust the agent number $n$ in MHSCA as 49, 144 and 256 for comparison, respectively.
Firstly, looking at the performance on both REC and RES tasks, our proposed MHSCA ($n$=256) achieves the overall best results. MHSCA ($n$=144) also obtains competitive performances. It is noteworthy that for visual feature maps with increasing spatial sizes, the FLOPs of our MHSCA never exceed 10M and demonstrate smaller growth compared to flash attention-based (Dao et al., 2022) MHCA and MHLCA. Regarding expanding channels, our MHSCA maintains a significantly lower parameter count compared to the other two fusion methods. Further, Figure 10 visualizes the last stage of PAN after fusing multi-scale attention maps under settings of MHSCA and MHCA, MHSCA correctly capture targets referred by prompt, also advanced in small target localization. Therefore, MHSCA achieves a well-balanced result between accuracy and complexity, serving as a lite and scalable cross-attention method for multi-scale maps.
Figure 10. Last-stage heatmaps of PAN before prediction heads.
Table 4. Comparison of REC on RefCOCOs datasets.
Models RefCOCO RefCOCO+ RefCOCOg
testA testB testA testB test-u
MAttNet 80.43 69.28 70.26 56.00 67.27
TransVG 82.72 78.35 70.70 56.94 67.73
RefTrans 85.53 76.31 75.58 61.91 69.10
MCN 82.29 74.98 72.86 57.31 66.01
MCN-MHSCA 83.06$\uparrow$ 75.22$\uparrow$ 73.54$\uparrow$ 58.67$\uparrow$ 66.92$\uparrow$
SeqTR 85.00 76.08 75.37 58.78 71.58
SeqTR-MHSCA 85.36$\uparrow$ 76.72$\uparrow$ 75.85$\uparrow$ 59.10$\uparrow$ 72.13$\uparrow$
Potamoi 84.98 78.10 75.90 58.55 70.55
Table 5. Comparison of RES on RefCOCOs datasets.
Models RefCOCO RefCOCO+ RefCOCOg
testA testB testA testB test-u
MAttNet 62.37 51.70 52.39 40.08 48.61
 VLT 68.29 62.73 59.20 49.36 56.65
MCN 64.20 59.71 54.99 44.69 49.40
MCN-MHSCA 64.51$\uparrow$ 59.94$\uparrow$ 55.30$\uparrow$ 45.32$\uparrow$ 49.76$\uparrow$
SeqTR 69.79 64.12 58.93 48.19 55.64
SeqTR-MHSCA 70.10$\uparrow$ 64.32$\uparrow$ 59.09$\uparrow$ 48.33$\uparrow$ 55.90$\uparrow$
Potamoi 68.71 63.02 58.70 48.69 54.12
Comparison of other settings. Table 6 presents ablation experiments of other settings in fusion-based Potamoi, including settings of input data, network and training. For radar features, target range and velocity are more vital for visual grounding. Besides, due to some samples containing prompt token lengths exceeding 30, and certain target-descriptive terms occurring towards the end of text, reducing token length resulted in information loss, leading to a decrease in accuracy. Moreover, when removing ARW and only with element-wise addition for vision-radar fusion, an apparent drop appears. When adopting multi-task learning, the performance of Potamoi is better than single-task learning, which indicates multi-task learning can promote performances of both tasks.
Table 6. Ablation experiments on other settings of Potamoi.
[…]
Baseline 70.1 44.8 53.0 70.81
Input Data
Radar map (R) 70.0 42.8 52.5 70.41
Radar map (R-V) 70.1 43.1 52.4 70.55
Radar map (R-P) 69.8 42.9 52.4 70.24
Radar map (V-P) 68.6 42.7 51.9 68.23
Token Len=30 69.8 43.2 52.0 70.31
Network Settings
No ARW 69.5 42.8 52.2 70.40
Training
Single-task 70.9 44.0 52.6 70.68
6. Conclusion
We propose the first visual grounding dataset of USV-based waterway perception, WaterVG, which includes fine-grained textual prompts, visual, and millimeter-wave radar data. Further, each sample’s text prompt contains references to multiple targets. Based on this, we design a multi-task visual grounding model, Potamoi, which can accurately performing REC and RES tasks simultaneously with the lowest power consumption among all models. This capability relies on our proposed ARW and highly efficient Multi-Head Slim Cross Attention (MHSCA) for fusion of textual and sensor features with a considerably low cost. Both Potamoi and MHSCA demonstrate competitive performance compared to counterparts.